Problem 4 Use Excel, to create a histogram for the data set. Calculate the cumulate distribution (CDF), remembering that it cannot exceed a value of "1", and plot it on the histogram. The histogram requires the creation of bins with appropriate labels and the frequency of occurrence in the data set. Properly document the plots.

14.81 7.92 10.50 22.59 5.62 8.01 6.53
10.62 5.11 6.28 9.71 3.13 3.99 7.50
8.68 11.69 17.81 8.70 5.38 7.43 14.27
7.52 4.09 8.71 18.17 5.10 9.66 6.44
13.13 25.14 6.39 6.89 9.29 10.46 8.88
9.95 7.80 5.28 5.42 6.69 7.94 8.56
6.11 7.59 10.84 5.60 25.61 11.59 20.48
12.95 11.63 18.44 14.60 15.48 8.96 11.76
10.55 7.48 11.76 6.67 23.19 5.11 10.24
12.51 10.03 7.33 19.18 9.15 6.40 17.09
9.43 7.78 11.69 13.60 12.24 8.79 9.33
5.50 15.01 10.89 7.44 9.16 12.50 14.81
8.78 10.78 10.36 9.63 13.28 7.85 9.07
8.05 17.98 8.68 5.39 10.25 16.51 4.99
8.98 23.00 8.46 13.14 14.69 9.80 21.97
7.46 13.90 10.29 5.65 9.61 7.02 13.76
10.88 7.66 6.74 6.80 13.05 9.64 20.54
17.18 10.05 7.03 5.97 12.92 6.61 17.09
17.74 7.72 6.65 7.13 10.68 10.77 6.72
13.54 8.11 17.39 8.93 24.03 10.80 7.83
Answer #1

How to draw Cumulative distributive frequency table:-

1) The data should not be in more than 1 column.Set all the data in 1 column.

2) After that,sort the data by using Excel function sort .

3) After that we use the formula that c2(column2) = (1/no of datapoints)+c1

In our example :-

C2 = 0.007142857143

C3 = (1/140)+0.007142857143 = 0.01428571429

similarly for all data points

Then we can plot the graph or any chart like histogram or scatter plot with lines.

Datapoints vs its frequency data points (count) (count) Frequency - - - - - - - - 29.00 30.00 11.00 12.00 data points 13.00 1

Graph with data points:-

=(1/140)+C2 B D E F G H I J K L Datapoints vs its frequency data points (count) (count) data points 3.13 0.007142857143 3.99

vs data points 0.75 0.5 6.67 0.1785714286 6.69 0.1857142857 6.72 0.1928571429 6.74 0.2 6.8 0.2071428571 6.89 0.2142857143 7.0

Here i provide google sheets link you can view in that file.

naveenvayilapalli/pdf-distribution in github,here you can download related cdf excel fiel or pdf file

If any doubts you have comment .

I hope you that you will like the answer.